OpenStudy (anonymous):

Write an equation in slope-intercept form for an equation of a line perpendicular to y = -3x +1 and that intersects at point (1, 2/3).

OpenStudy (anonymous):

POP, POP!

OpenStudy (anonymous):

first for perpendicular lines you switch the sign of the -3 so its positive

OpenStudy (anonymous):

y-y1=m(x-x1) y-2/3=3(x-1) and solve

OpenStudy (anonymous):

distribute 3 to the x and -1 then add 2/3